Run Plan Nov. 21 - Dec 6 2005 Last altered Wednesday Nov. 30 /------------------------------------------------------------------ Monday Nov 21 Find settings for pure Al stops. -achieved by maximising the number of "last hits" in PC6 compared with PC5. -idea is to move the stopping distribution so that we minimise the number of muons stopping in gas upstream of the target, leaving muons decaying in the target (which we keep) and muons decaying downstream (which we can remove). -meant to study the polarisation of the muons in the target without the depolarising influence of muons stopping in gas. when a satisfactory setting is achieved set for stable running of this data set until maintenance **Historical note** a High rate test was also performed while searching for an appropriate gas degrader setting. /------------------------------------------------------------------ Tuesday Nov 22 continue running pure Al stops data set /------------------------------------------------------------------ Wednesday Nov 23 maintenance starts at 8:30. End Al stops data install TEC at this time (with 49 hole collimator) ("4 hour" maintenance) start with collimator data using Nov. 18 spray tune /----------------------------------------------------------------- Thursday Nov 24 continue TEC collimator data /----------------------------------------------------------------- Friday Nov 25 end TEC collimator data ~4:00 pm remove 49 hole collimator START NOMINAL DATA SET -remove TEC -use the settings found on the nominal settings page /----------------------------------------------------------------- Saturday Nov 26 continue nominal data set /---------------------------------------------------------------- Sunday Nov 27 end nominal data set Sunday morning start B2 + 6G data set /---------------------------------------------------------------- Monday Nov 28 continue B2 + 6G data set /---------------------------------------------------------------- Tuesday Nov 29 B2 + 6G with the beginning of Maintenance day /--------------------------------------------------------------- Wednesday Nov 30 Insert TEC in nominal configuration during Maintenance When beam returns, set up far upstream stops data set Far upstream stops set - compare data in this geometry to far upstream stops in the old geometry - details of setup to be filled later: reference runs 18693-18700 are archived on Westgrid /--------------------------------------------------------------- Thursday Dec 1 Take some characterisation runs of the beam at "nominal" and "B2+6G" settings - approximatly 2 runs for each setting - this gives us a charactorisation of the for the previous 2 data sets - produces a baseline for the comparison of the reversed TEC configuration Reverse the X and Y modules in the planes - ideally this should tell us whether the observed Y angle is postion dependent or not. - initially introduce TEC in normal configuration. Take a few runs to provide a baseline to compare with the reversed case. - reverse the planes. Take a few runs to compare the operation in these two cases. steering tests - presumably set up during the maintenance day - attempt to use the power sources we have to steer the beam overnight (or after all the tests made during the day are completed) Far upstream stops set continued from overnight Wednesday /--------------------------------------------------------------- Friday Dec 2 end far upstream stops set. (should be in excess of 100 runs after the two nights) t0 run with field on, MLT trigger. Note that this requires a 120 MeV/c focused (non-spray) pion tune (to increase the rate) Ramp down to 0T alignment data /--------------------------------------------------------------- Saturday Dec 3 -continue alignment data -t0 data- use MLT, PU, DS trigger scintillators (4 runs for each) /---------------------------------------------------------------- Sunday Dec 4 /--------------------------------------------------------------- Monday Dec 5 -remove TEC and beam package -insert MTEC -quads off TEC beam measurements to see if TEC/beam relationship consistant with Aug. 2005=>TEC postion -Install Big collimators -run with this collimator (to check absolute postion of beam in spectrometer?) /--------------------------------------------------------------- Tuesday -beam positron characterisation run -big collimators/ TEC out -large DS paddle scintillator -"one or two runs should be sufficient"